# Reality conditions for the KdV equation and quasi-periodic solutions in finite phase spaces

### v2

@@ 3.1. The phase space of KdV hierarchy
$\mathcal{M}_{N}=\bigg{\{}\Psi=\sum_{\ell=-1}^{2N+1}\sum_{a=1,2,3}\psi_{a,\ell}Z_{a,\ell}^{\ast}\bigg{\}},$
@@
(9c) $\Psi(z)=\begin{pmatrix}\alpha(z)&\beta(z)\\ \gamma(z)&-\alpha(z)\end{pmatrix},$
(9d) $\alpha(z)=\sum_{m=0}^{N}\alpha_{2i}z^{m},\quad\beta(z)=\sum_{m=0}^{N+1}\beta_{2i-1}z^{m-1},\quad\gamma(z)=\sum_{m=0}^{N+1}\gamma_{2m-1}z^{m}.$
@@
Initial points $\Phi$ can be taken from the Weyl chamber of $\widetilde{G}_{-}$ in $\mathcal{M}_{N}$. The Weyl chamber is spanned by $H_{2m}^{\ast}$, $m=0$, …, $N$. Thus, initial points $\Phi$ are given by diagonal matrices.
@@
(10b) $W_{i,j}^{a,b}=\langle\Psi,[Z_{a,-i-2},Z_{a,-j-2}]\rangle.$
@@
In terms of the dynamic variables the symplectic structure (3.1) is defined by
@@
In $\mathcal{M}_{N}$ with $\Psi$ of the form (3.1), coadjoint orbits of $\widetilde{G}_{-}$ serve as finite phase spaces for the hierarchy of the $+$mKdV equation in the case of $\mathfrak{g}=\mathfrak{su}(2)$, and --mKdV in the case of $\mathfrak{g}=\mathfrak{sl}(2,\operatorname{\mathbb{R}})$. At the same time, coadjoint orbits of $\widetilde{G}_{+}=\exp(\widetilde{\mathfrak{g}}_{+})$ serve as phase spaces for the hierarchies of the $\sin$-Gordon or $\sinh$-Gordon equation, respectively.
@@
(12) $\mathfrak{g}_{-1}=\operatorname{span}\{X_{-1}+Y_{-1}\},\qquad\beta_{-1}=0.$
@@
$\textsf{W}=\begin{pmatrix}0&\overset{\circ}{\textsf{w}}_{1}&\overset{\circ}{\textsf{w}}_{2}&\dots&\overset{\circ}{\textsf{w}}_{N-1}&\overset{\circ}{\textsf{w}}_{N}\\ -\overset{\circ}{\textsf{w}}{}_{1}^{t}&\textsf{w}_{2}&\textsf{w}_{3}&\dots&\textsf{w}_{N}&\textsf{w}_{N+1}\\ -\overset{\circ}{\textsf{w}}{}_{2}^{t}&\textsf{w}_{3}&\textsf{w}_{4}&\dots&\textsf{w}_{N+1}&0\\ \vdots&\vdots&\vdots&\iddots&0&0\\ -\overset{\circ}{\textsf{w}}_{N-1}^{t}&\textsf{w}_{N}&\textsf{w}_{N+1}&\iddots&0&0\\ -\overset{\circ}{\textsf{w}}_{N}^{t}&\textsf{w}_{N+1}&0&\dots&0&0\end{pmatrix},$
@@ 3.3. KdV equation
On $\mathcal{M}_{N}^{\circ}$ we consider two hamiltonians: $h_{N-1}$ gives rise to a stationary flow with parameter $x$, and $h_{N-2}$ gives rise to an evolutionary flow with parameter $t$:
(18) $\frac{\mathrm{d}\psi_{a,\ell}}{\mathrm{d}x}=\{h_{N-1},\psi_{a,\ell}\},\qquad\frac{\mathrm{d}\psi_{a,\ell}}{\mathrm{d}t}=\{h_{N-2},\psi_{a,\ell}\}.$
@@
$\frac{\mathrm{d}\gamma_{-1}}{\mathrm{d}x}=-2\alpha_{0}(\beta_{2N-1}-\gamma_{2N-1})-2\mathrm{a}\gamma_{-1},$
$\frac{\mathrm{d}\alpha_{2m}}{\mathrm{d}x}=-\mathrm{b}(\beta_{2m-1}-\gamma_{2m-1})+\beta_{2m+1}(\beta_{2N-1}-\gamma_{2N-1}),$
$\frac{\mathrm{d}\beta_{2m+1}}{\mathrm{d}x}=-2\mathrm{b}\alpha_{2m}+2\mathrm{a}\beta_{2m+1},$
$\frac{\mathrm{d}\gamma_{2m+1}}{\mathrm{d}x}=2\mathrm{b}\alpha_{2m}-2\alpha_{2m+2}(\beta_{2N-1}-\gamma_{2N-1})-2\mathrm{a}\gamma_{2m+1},\quad m=0,\dots,N-1.$
@@
$\frac{\mathrm{d}\beta_{2N-1}}{\mathrm{d}t}=-2\mathrm{b}\alpha_{2N-4}+2\mathrm{a}\beta_{2N-3}.$
@@
(19) $\frac{\mathrm{d}\beta_{2N-1}}{\mathrm{d}t}=\frac{\mathrm{d}\beta_{2N-3}}{\mathrm{d}x},$
@@
$\frac{\mathrm{d}\alpha_{2N-2}}{\mathrm{d}x}=-\mathrm{b}(\beta_{2N-3}-\gamma_{2N-3})+\beta_{2N-1}(\beta_{2N-1}-\gamma_{2N-1}),$
@@
(20) $2\mathrm{b}\beta_{2N-3}=-\frac{\mathrm{d}\alpha_{2N-2}}{\mathrm{d}x}+\beta_{2N-1}(\beta_{2N-1}-2\gamma_{2N-1})-2\mathrm{a}\alpha_{2N-2}+c_{2N-1}.$
@@
$\frac{\mathrm{d}\beta_{2N-1}}{\mathrm{d}x}=-2\mathrm{b}\alpha_{2N-2}+2\mathrm{a}\beta_{2N-1},$
@@
$\alpha_{2N-2}=-\frac{1}{2\mathrm{b}}\frac{\mathrm{d}\beta_{2N-1}}{\mathrm{d}x}+\frac{\mathrm{a}}{\mathrm{b}}\beta_{2N-1},$
@@
(21) $\beta_{2N-3}=\frac{1}{4\mathrm{b}^{2}}\Big{(}\frac{\mathrm{d}^{2}\beta}{\mathrm{d}x^{2}}+6\mathrm{b}\beta^{2}-4c_{2N}\beta+2c_{2N-1}\mathrm{b}\Big{)},$
where $\beta$ stands for $\beta_{2N-1}$. Finally, differentiating (21) with respect to $x$ and substituting into (19), we obtain
(22) $\frac{\mathrm{d}\beta}{\mathrm{d}t}=\frac{1}{4\mathrm{b}^{2}}\Big{(}\frac{\mathrm{d}^{3}\beta}{\mathrm{d}x^{3}}+12\mathrm{b}\beta\frac{\mathrm{d}\beta}{\mathrm{d}x}-4c_{2N}\frac{\mathrm{d}\beta}{\mathrm{d}x}\Big{)},$
@@
(23a) $\frac{\mathrm{d}\alpha_{0}}{\mathrm{d}x}=\beta_{1}^{2}+\mathrm{b}\gamma_{-1}-\beta_{1}\gamma_{1},$
(23b) $\frac{\mathrm{d}\beta_{1}}{\mathrm{d}x}=-2\mathrm{b}\alpha_{0}+2\mathrm{a}\beta_{1}.$
@@
(24) $\frac{\mathrm{d}^{2}\beta_{1}}{\mathrm{d}x^{2}}+6\mathrm{b}\beta_{1}^{2}-4c_{2}\beta_{1}=-2\mathrm{b}c_{1},$
@@ 3.4. Higher KdV equations
$\frac{\mathrm{d}\beta_{2N-1}}{\mathrm{d}\tau_{m-1}}=\frac{\mathrm{d}\beta_{2m-1}}{\mathrm{d}x},\quad m=1,\,\dots,\,N-2,$
@@
$\frac{\mathrm{d}\beta}{\mathrm{d}\tau_{N-3}}=\frac{1}{16\mathrm{b}^{4}}\bigg{(}\frac{\mathrm{d}^{5}\beta}{\mathrm{d}x^{5}}+20\mathrm{b}\Big{(}\beta\frac{\mathrm{d}^{3}\beta}{\mathrm{d}x^{3}}+2\frac{\mathrm{d}\beta}{\mathrm{d}x}\frac{\mathrm{d}^{2}\beta}{\mathrm{d}x^{2}}\Big{)}+120\mathrm{b}^{2}\beta^{2}\frac{\mathrm{d}\beta}{\mathrm{d}x}\\ +8\mathrm{b}^{2}c_{2N-1}\frac{\mathrm{d}\beta}{\mathrm{d}x}\bigg{)},$
which coincides with [22, Eq. (8’’), p. 244].
@@ 3.5. Zero curvature representation
$\frac{\mathrm{d}\Psi}{\partial x}=[\Psi,\nabla h_{N-1}],\qquad\frac{\mathrm{d}\Psi}{\partial t}=[\Psi,\nabla h_{N-2}],$
@@
$\nabla h_{n}=\sum_{i=-1}^{2N+1}\sum_{a=1,2,3}\frac{\partial h_{n}}{\partial\psi_{a,i}}Z_{a,-i-2}.$
The matrix gradient of each flow has a complementary matrix $A$, such that
$[\Psi,\nabla h_{n}]=[\Psi,A].$
@@
(25) $\frac{\mathrm{d}\Psi}{\partial x}=[\Psi,A_{\text{st}}],\qquad\frac{\mathrm{d}\Psi}{\partial t}=[\Psi,A_{\text{ev}}],$
$\begin{split}&A_{\text{st}}=-\begin{pmatrix}\mathrm{a}&\mathrm{b}\\ \mathrm{b}z+\gamma_{2N-1}-\beta_{2N-1}&-\mathrm{a}\end{pmatrix},\\ &A_{\text{ev}}=-\begin{pmatrix}\mathrm{a}z+\alpha_{2N-2}&\mathrm{b}z+\beta_{2N-1}\\ \mathrm{b}z^{2}+\gamma_{2N-1}z+\gamma_{2N-3}-\beta_{2N-3}&-(\mathrm{a}z+\alpha_{2N-2})\end{pmatrix}.\end{split}$
@@
$\frac{\mathrm{d}A_{\text{st}}}{\partial t}-\frac{\mathrm{d}A_{\text{ev}}}{\partial x}=[A_{\text{st}},A_{\text{ev}}].$
@@ 3.6. Summary
The loop algebra $\widetilde{\mathfrak{g}}=\mathfrak{sl}(2,\operatorname{\mathbb{R}})\otimes\mathcal{P}(z,z^{-1})$ with the principal grading is associated with the KdV hierarchy. Let $\mathcal{M}^{\circ}_{N}$ be the manifold $\widetilde{\mathfrak{g}}^{\ast}_{-}/\big{(}\sum_{\ell\geqslant 2N+2}\mathfrak{g}_{\ell}\big{)}$ with the hamiltonian reduction (12), $\dim\mathcal{M}^{\circ}_{N}=3N+1$, $N\,{\in}\,\operatorname{\mathbb{N}}$. Evidently, $\mathcal{M}^{\circ}_{1}\subset\mathcal{M}^{\circ}_{2}\subset\cdots\subset\mathcal{M}^{\circ}_{N}\subset\mathcal{M}^{\circ}_{N+1}\subset\cdots$. Each manifold $\mathcal{M}^{\circ}_{N}$ is equipped with the symplectic structure (13). Under the action of the loop group $\widetilde{G}_{-}=\exp(\widetilde{\mathfrak{g}}_{-})$ a manifold $\mathcal{M}^{\circ}_{N}$ splits into orbits $\mathcal{O}$, each generated by a point from the Weyl chamber. On the other hand, such an orbit is defined by the system of $N+1$ constraints (17). Each orbit serves as a phase space of dimension $2N$ for a hamiltonian system integrable in the Liouville sence.
On orbits within $\mathcal{M}^{\circ}_{N}$, $N\geqslant 2$, there exist two hamiltonians whose flows give rise to the KdV equation (22). We call these flows stationary and evolutionary with parameters $x$ and $t$, correspondingly. On orbits in $\mathcal{M}^{\circ}_{1}$ there exists a stationary flow only. If $N>2$, higher KdV equations arise. One can use the remaining hamiltonians to generate evolutionary flows.

### v4

@@ 3.1. The phase space of KdV hierarchy
$\mathcal{M}_{N}=\bigg{\{}\Psi=\sum_{\ell=-1}^{2N+1}\sum_{a=1,2,3}\psi_{a,\ell}\textsf{Z}_{a,\ell}^{\ast}\bigg{\}},$
@@
(13a) $\Psi(z)=\begin{pmatrix}\alpha(z)&\beta(z)\\ \gamma(z)&-\alpha(z)\end{pmatrix},$
(13b) $\alpha(z)=\sum_{m=0}^{N}\alpha_{2m}z^{m},\quad\beta(z)=\sum_{m=0}^{N+1}\beta_{2m-1}z^{m-1},\quad\gamma(z)=\sum_{m=0}^{N+1}\gamma_{2m-1}z^{m}.$
@@
Initial points $\Phi$ can be taken from the Weyl chamber of $\widetilde{G}_{-}$ in $\mathcal{M}_{N}$. The Weyl chamber is spanned by $\textsf{H}_{2m}^{\ast}$, $m=0$, …, $N$. Thus, initial points $\Phi$ are given by diagonal matrices.
@@
(14b) $W_{i,j}^{a,b}=\langle\Psi,[\textsf{Z}_{a,-i-2},\textsf{Z}_{a,-j-2}]\rangle.$
@@
In terms of the dynamic variables the symplectic structure (14) is defined by
@@
In $\mathcal{M}_{N}$ with $\Psi$ of the form (13), coadjoint orbits of $\widetilde{G}_{-}$ serve as finite phase spaces for the hierarchy of the $+$mKdV equation in the case of $\mathfrak{g}=\mathfrak{su}(2)$, and --mKdV in the case of $\mathfrak{g}=\mathfrak{sl}(2,\operatorname{\mathbb{R}})$. At the same time, coadjoint orbits of $\widetilde{G}_{+}=\exp(\widetilde{\mathfrak{g}}_{+})$ serve as phase spaces for the hierarchies of the $\sin$-Gordon or $\sinh$-Gordon equation, respectively.
@@
(16) $\mathfrak{g}_{-1}=\operatorname{span}\{\textsf{X}_{-1}+\textsf{Y}_{-1}\},\qquad\beta_{-1}=0.$
@@
$\textsf{W}=\begin{pmatrix}0&\overset{\circ}{\textsf{w}}_{1}&\overset{\circ}{\textsf{w}}_{2}&\dots&\overset{\circ}{\textsf{w}}_{N-1}&\overset{\circ}{\textsf{w}}_{N}\\ -\overset{\circ}{\textsf{w}}{}_{1}^{t}&\textsf{w}_{2}&\textsf{w}_{3}&\dots&\textsf{w}_{N}&\textsf{w}_{N+1}\\ -\overset{\circ}{\textsf{w}}{}_{2}^{t}&\textsf{w}_{3}&\textsf{w}_{4}&\dots&\textsf{w}_{N+1}&0\\ \vdots&\vdots&\vdots&\iddots&0&0\\ -\overset{\circ}{\textsf{w}}_{N-1}^{t}&\textsf{w}_{N}&\textsf{w}_{N+1}&\iddots&\vdots&\vdots\\ -\overset{\circ}{\textsf{w}}_{N}^{t}&\textsf{w}_{N+1}&0&\dots&0&0\end{pmatrix},$
@@ 3.3. KdV equation
On $\mathcal{M}_{N}^{\circ}$ we consider two hamiltonians: $h_{N-1}$ gives rise to a stationary flow with parameter $\mathrm{x}$, and $h_{N-2}$ gives rise to an evolutionary flow with parameter $\mathrm{t}$:
(22) $\frac{\mathrm{d}\psi_{a,\ell}}{\mathrm{d}\mathrm{x}}=\{h_{N-1},\psi_{a,\ell}\},\qquad\frac{\mathrm{d}\psi_{a,\ell}}{\mathrm{d}\mathrm{t}}=\{h_{N-2},\psi_{a,\ell}\}.$
@@
$\frac{\mathrm{d}\gamma_{-1}}{\mathrm{d}\mathrm{x}}=-2\alpha_{0}(\beta_{2N-1}-\gamma_{2N-1})-2\mathrm{a}\gamma_{-1},$
$\frac{\mathrm{d}\alpha_{2m}}{\mathrm{d}\mathrm{x}}=-\mathrm{b}(\beta_{2m-1}-\gamma_{2m-1})+\beta_{2m+1}(\beta_{2N-1}-\gamma_{2N-1}),$
$\frac{\mathrm{d}\beta_{2m+1}}{\mathrm{d}\mathrm{x}}=-2\mathrm{b}\alpha_{2m}+2\mathrm{a}\beta_{2m+1},$
$\frac{\mathrm{d}\gamma_{2m+1}}{\mathrm{d}\mathrm{x}}=2\mathrm{b}\alpha_{2m}-2\alpha_{2m+2}(\beta_{2N-1}-\gamma_{2N-1})-2\mathrm{a}\gamma_{2m+1},\quad m=0,\dots,N-1.$
@@
$\frac{\mathrm{d}\beta_{2N-1}}{\mathrm{d}\mathrm{t}}=-2\mathrm{b}\alpha_{2N-4}+2\mathrm{a}\beta_{2N-3}.$
@@
(23) $\frac{\mathrm{d}\beta_{2N-1}}{\mathrm{d}\mathrm{t}}=\frac{\mathrm{d}\beta_{2N-3}}{\mathrm{d}\mathrm{x}},$
@@
$\frac{\mathrm{d}\alpha_{2N-2}}{\mathrm{d}\mathrm{x}}=-\mathrm{b}(\beta_{2N-3}-\gamma_{2N-3})+\beta_{2N-1}(\beta_{2N-1}-\gamma_{2N-1}),$
@@
(24) $2\mathrm{b}\beta_{2N-3}=-\frac{\mathrm{d}\alpha_{2N-2}}{\mathrm{d}\mathrm{x}}+\beta_{2N-1}(\beta_{2N-1}-2\gamma_{2N-1})-2\mathrm{a}\alpha_{2N-2}+c_{2N-1}.$
@@
$\frac{\mathrm{d}\beta_{2N-1}}{\mathrm{d}\mathrm{x}}=-2\mathrm{b}\alpha_{2N-2}+2\mathrm{a}\beta_{2N-1},$
@@
$\alpha_{2N-2}=-\frac{1}{2\mathrm{b}}\frac{\mathrm{d}\beta_{2N-1}}{\mathrm{d}\mathrm{x}}+\frac{\mathrm{a}}{\mathrm{b}}\beta_{2N-1},$
@@
(25) $\beta_{2N-3}=\frac{1}{4\mathrm{b}^{2}}\Big{(}\frac{\mathrm{d}^{2}\beta}{\mathrm{d}\mathrm{x}^{2}}+6\mathrm{b}\beta^{2}-4c_{2N}\beta+2c_{2N-1}\mathrm{b}\Big{)},$
where $\beta$ stands for $\beta_{2N-1}$. Finally, differentiating (25) with respect to $\mathrm{x}$ and substituting into (23), we obtain
(26) $\frac{\mathrm{d}\beta}{\mathrm{d}\mathrm{t}}=\frac{1}{4\mathrm{b}^{2}}\Big{(}\frac{\mathrm{d}^{3}\beta}{\mathrm{d}\mathrm{x}^{3}}+12\mathrm{b}\beta\frac{\mathrm{d}\beta}{\mathrm{d}\mathrm{x}}-4c_{2N}\frac{\mathrm{d}\beta}{\mathrm{d}\mathrm{x}}\Big{)},$
@@
(27a) $\frac{\mathrm{d}\alpha_{0}}{\mathrm{d}\mathrm{x}}=\beta_{1}^{2}+\mathrm{b}\gamma_{-1}-\beta_{1}\gamma_{1},$
(27b) $\frac{\mathrm{d}\beta_{1}}{\mathrm{d}\mathrm{x}}=-2\mathrm{b}\alpha_{0}+2\mathrm{a}\beta_{1}.$
@@
(28) $\frac{\mathrm{d}^{2}\beta_{1}}{\mathrm{d}\mathrm{x}^{2}}+6\mathrm{b}\beta_{1}^{2}-4c_{2}\beta_{1}=-2\mathrm{b}c_{1},$
@@ 3.4. Higher KdV equations
$\frac{\mathrm{d}\beta_{2N-1}}{\mathrm{d}\tau_{m-1}}=\frac{\mathrm{d}\beta_{2m-1}}{\mathrm{d}\mathrm{x}},\quad m=1,\,\dots,\,N-2,$
@@
$\frac{\mathrm{d}\beta}{\mathrm{d}\tau_{N-3}}=\frac{1}{16\mathrm{b}^{4}}\bigg{(}\frac{\mathrm{d}^{5}\beta}{\mathrm{d}\mathrm{x}^{5}}+20\mathrm{b}\Big{(}\beta\frac{\mathrm{d}^{3}\beta}{\mathrm{d}\mathrm{x}^{3}}+2\frac{\mathrm{d}\beta}{\mathrm{d}\mathrm{x}}\frac{\mathrm{d}^{2}\beta}{\mathrm{d}\mathrm{x}^{2}}\Big{)}+120\mathrm{b}^{2}\beta^{2}\frac{\mathrm{d}\beta}{\mathrm{d}\mathrm{x}}\\ +8\mathrm{b}^{2}c_{2N-1}\frac{\mathrm{d}\beta}{\mathrm{d}\mathrm{x}}\bigg{)},$
which coincides with [33, Eq. (8’’), p. 244].
@@ 3.5. Zero curvature representation
$\frac{\mathrm{d}\Psi}{\partial\mathrm{x}}=[\Psi,\nabla h_{N-1}],\qquad\frac{\mathrm{d}\Psi}{\partial\mathrm{t}}=[\Psi,\nabla h_{N-2}],$
@@
$\nabla h_{n}=\sum_{i=-1}^{2N+1}\sum_{a=1,2,3}\frac{\partial h_{n}}{\partial\psi_{a,i}}\textsf{Z}_{a,-i-2}.$
The matrix gradient of each flow has a complementary matrix $\mathrm{A}$, such that
$[\Psi,\nabla h_{n}]=[\Psi,\mathrm{A}].$
@@
(29) $\frac{\mathrm{d}\Psi}{\partial\mathrm{x}}=[\Psi,\mathrm{A}_{\text{st}}],\qquad\frac{\mathrm{d}\Psi}{\partial\mathrm{t}}=[\Psi,\mathrm{A}_{\text{ev}}],$
$\begin{split}&\mathrm{A}_{\text{st}}=-\begin{pmatrix}\mathrm{a}&\mathrm{b}\\ \mathrm{b}z+\gamma_{2N-1}-\beta_{2N-1}&-\mathrm{a}\end{pmatrix},\\ &\mathrm{A}_{\text{ev}}=-\begin{pmatrix}\mathrm{a}z+\alpha_{2N-2}&\mathrm{b}z+\beta_{2N-1}\\ \mathrm{b}z^{2}+\gamma_{2N-1}z+\gamma_{2N-3}-\beta_{2N-3}&-(\mathrm{a}z+\alpha_{2N-2})\end{pmatrix}.\end{split}$
@@
$\frac{\mathrm{d}\mathrm{A}_{\text{st}}}{\partial\mathrm{t}}-\frac{\mathrm{d}\mathrm{A}_{\text{ev}}}{\partial\mathrm{x}}=[\mathrm{A}_{\text{st}},\mathrm{A}_{\text{ev}}].$
@@ 3.6. Summary
The affine algebra $\widetilde{\mathfrak{g}}=\mathfrak{sl}(2,\operatorname{\mathbb{R}})\otimes\mathcal{P}(z,z^{-1})$ with the principal grading is associated with the KdV hierarchy. Let $\mathcal{M}^{\circ}_{N}$ be the manifold $\widetilde{\mathfrak{g}}^{\ast}_{-}/\big{(}\sum_{\ell\geqslant 2N+2}\mathfrak{g}_{\ell}\big{)}$ with the hamiltonian reduction (16), $\dim\mathcal{M}^{\circ}_{N}=3N+1$, $N\,{\in}\,\operatorname{\mathbb{N}}$. Evidently, $\mathcal{M}^{\circ}_{1}\subset\mathcal{M}^{\circ}_{2}\subset\cdots\subset\mathcal{M}^{\circ}_{N}\subset\mathcal{M}^{\circ}_{N+1}\subset\cdots$. Each manifold $\mathcal{M}^{\circ}_{N}$ is equipped with the symplectic structure (17). Under the action of the loop group $\widetilde{G}_{-}=\exp(\widetilde{\mathfrak{g}}_{-})$ a manifold $\mathcal{M}^{\circ}_{N}$ splits into orbits $\mathcal{O}$, each generated by a point from the Weyl chamber. On the other hand, such an orbit is defined by the system of $N+1$ constraints (21). Each orbit serves as a phase space of dimension $2N$ for a hamiltonian system integrable in the Liouville sence.
On orbits within $\mathcal{M}^{\circ}_{N}$, $N\geqslant 2$, there exist two hamiltonians whose flows give rise to the KdV equation (26). We call these flows stationary and evolutionary with parameters $\mathrm{x}$ and $\mathrm{t}$, correspondingly. On orbits in $\mathcal{M}^{\circ}_{1}$ there exists a stationary flow only. If $N>2$, higher KdV equations arise. One can use the remaining hamiltonians to generate evolutionary flows.